### v1

# “Whose Side Are You On?” Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection

Abstract
The rapid growth of social media platforms has led to concerns about radicalization, filter bubbles, and content bias. Existing approaches to classifying ideology are limited in that they require extensive human effort, the labeling of large datasets, and are not able to adapt to evolving ideological contexts. This paper explores the potential of Large Language Models (LLMs) for classifying the political ideology of online content in the context of the two-party US political spectrum through in-context learning (ICL). Our extensive experiments involving demonstration selection in label-balanced fashion, conducted on three datasets comprising news articles and YouTube videos, reveal that our approach significantly outperforms zero-shot and traditional supervised methods. Additionally, we evaluate the influence of metadata (e.g., content source and descriptions) on ideological classification and discuss its implications. Finally, we show how providing the source for political and non-political content influences the LLM’s classification.
“Whose Side Are You On?”
Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection
Muhammad Haroon University of California, Davis mharoon@ucdavis.edu Magdalena Wojcieszak University of California, Davis mwojcieszak@ucdavis.edu Anshuman Chhabra University of South Florida anshumanc@usf.edu
1 Introduction
Social media platforms are one of the leading sources of information for the majority of Americans and have grown in centrality over the past few years Barthel et al. (2020). Observers worry that the content recommended and seen on social media platforms can radicalize users, cause extremism Tufekci (2018); Roose (2019), and facilitate the formation of filter bubbles that restrict exposure to diverse viewpoints Pariser (2011).
However, empirical evidence for these worries is inconsistent Haroon et al. (2023); Hosseinmardi et al. (2021); Chen et al. (2023); Nyhan et al. (2023) in part because scholars use different methodologies and approaches (audits versus observational data) and in part because we lack of robust (i.e., highly accurate across domains) and scalable (i.e., requiring minimal human supervision) methods for estimating political ideology of the content users consume.111An orthogonal problem is estimating ideology of users themselves Barberá (2015) but is outside our paper’s scope. Systematic assessments of whether (and the extent to which) social media platforms indeed expose users to ideologically biased or politically radical content is not possible without robust and scalable political ideology estimation methods.
Prior approaches to ideology estimation have utilized manual expert annotations Otero (2019), supervised classification models Baly et al. (2020), and social media sharing patterns Le et al. (2017); Barberá (2015); Lai et al. (2024); Wojcieszak et al. (2023); Eady et al. (2020). Manual annotations are highly accurate but resource-intensive and difficult to scale, especially given the growing volume of online content. Supervised classification models, while offering some automation, face challenges such as a reliance on extensive labeled data and difficulty adapting to dynamic ideological contexts. Social media-based methods estimate ideology by analyzing content sharing patterns on platforms such as Twitter or Reddit, but these approaches are hindered by data access limitations Murtfeldt et al. (2024), reliance on widely shared content, and being agnostic to the content itself. These limitations highlight the need for scalable, adaptable, and robust solutions to ideology estimation.
In this paper, we show that Large Language Models (LLMs) can be applied to ideology estimation of text in the context of the two-party US political spectrum (democrat/liberal and republican/conservative) and can address the aforementioned challenges of scalability and robustness. More specifically, pre-trained LLMs can learn new classification tasks in-context at test-time Hardt and Sun , which enables a few-shot learning paradigm that only requires a few annotated samples. Furthermore, LLMs can handle complex language subtleties necessary for accurate/robust ideology estimation. They can also adapt to new and evolving information by simple adjustments of the provided prompt. Thus, LLMs can serve as a compelling choice for robust and scalable ideology estimation across diverse text domains and platforms.
Indeed, LLMs have recently been applied to this task Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023). These past applications have primarily focused on zero-shot approaches, utilizing pre-trained LLMs without employing more nuanced few-shot in-context learning (ICL) strategies, as we detail below. Our work advances these prior studies by proposing a modified version of coverage-based ICL that selects demonstrations in a class-balanced manner, significantly improving performance over simple zero-shot prompting and coverage-based ICL. Moreover, whereas existing works primarily classify political texts with explicit ideology (e.g., tweets from senators), we demonstrate the effectiveness of our approach across diverse content categories, including political news, political non-news, and non-political news content.
In sum, we make the following contributions:
We conduct an extensive study on the capabilities of various LLMs (Llama2-13B, Mistral-7B, and GPT-4o) in estimating the ideology of text. Through rigorous experiments on three distinct datasets, we demonstrate LLMs’ accuracy, adaptability, and robustness in capturing ideological cues across multiple domains.
We propose a demonstration selection algorithm that utilizes coverage-based ICL and class-balanced selection to achieve state-of-the-art LLM performance on ideology estimation. Our approach improves upon zero-shot prompting used in past work Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023).
We analyze the effect of additional metadata, such as content source and description, on predictions, especially for neutral or non-political content. Our results show that metadata can significantly influence LLM predictions, at times aligning the predicted ideology with that of the source or introducing bias when source and content ideologies diverge. Our work provides critical insights into the role of LLMs for nuanced ideological analysis.
2 Related Work
Prior approaches to ideology estimation have generally used (1) manual expert annotations Otero (2019), (2) content analysis through supervised classification models Baly et al. (2020), and (3) social media sharing approaches Le et al. (2017); Barberá (2015); Lai et al. (2024); Wojcieszak et al. (2023); Eady et al. (2020).
Expert labeling. Several efforts have been taken by independent parties to identify the political ideology of online news articles and sources such as Ad Fontes Otero (2019), AllSides222https://allsides.com, Media Bias/Fact Check333https://mediabiasfactcheck.com, etc through a manual annotation process involving experts. Whereas manual annotation is accurate and robust (albeit depending on the number and quality of experts employed), it is highly resource-intensive and not scalable. Expert annotation becomes especially challenging due to the ever-increasing volume of social media content.
Supervised models. In turn, supervised ideology classification models can provide some form of automation. The advent of deep learning techniques revolutionized this field by enabling more sophisticated models capable of capturing complex linguistic patterns. Iyyer et al. (2014) introduced deep neural networks, such as LSTMs and CNNs, which outperformed traditional models in detecting ideology from unstructured text data. In a similar vein, Baly et al. (2020) used a BERT model trained on political news articles to show improvements over existing state-of-the-art supervised learning approaches. However, they are limited by a need for large volumes of human-labeled data and have difficulty adapting to dynamic and ever evolving ideological contexts (e.g., cross-platform estimation, emerging news events, changes in the political environment).
Social media sharing approaches. Social media sharing ideology estimation approaches utilize data from external platforms such as Twitter Le et al. (2017); Barberá (2015); Wojcieszak et al. (2023); Eady et al. (2020) or Reddit Lai et al. (2024) to assess how a given piece of content is shared among users. By identifying left- and right-leaning elites (e.g., politicians, news media) those users follow, those approaches first estimate the ideology of the users, and then estimate the ideology of the shared content by aggregating the ideology of the sharing users. These approaches are limited in several ways. First, as it relies on third-party platforms (such as Twitter/X), changes to APIs and data acquisition systems can render this approach untenable. For instance, since Twitter transitioned to X in 2023, it has become challenging for researchers to collect sharing data at scale without incurring significant costs Murtfeldt et al. (2024); Poudel and Weninger (2024). Second, to accurately estimate the ideology of content (e.g., a YouTube video), the content needs to be shared extensively, which might not happen for niche videos with small viewership. Finally, this approach is agnostic to the nature of the content itself and only uses sharing patterns.
Using LLMs. LLMs have achieved state-of-the-art performance for several tasks across multiple language domains (e.g., legal transcripts Colombo et al. (2024) and medical data Yang et al. (2022)) and contexts (e.g., identifying misinformation Hoes et al. (2023), hate speech Huang et al. (2023), sentiment Heseltine et al. (2024), or relevance, topics, and frames in tweets and news articles Gilardi et al. (2023).
Remark. Prior works have utilized LLMs to classify the ideology of text Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023). As we show in Table 1, our work differs from these in several key ways. First, all these approaches are zero-shot methods that only utilize the pre-trained LLMs as is, and do not employ optimal and efficient few-shot ICL approaches to improve accuracy and adapt LLMs to multiple domains. We propose a modified version of coverage-based ICL that selects demonstrations in a class-balanced manner, and improves performance over simple zero-shot prompting and coverage-based ICL. Second, these methods focus on classifying texts from sources with clearly distinct ideology (e.g. tweets from liberal or conservative politicians). We focus on content with differential levels of ideological signal to demonstrate that our LLM-based approach can estimate ideology of political news, political content but not news, and non-political news content as well. We also analyze the potential limitations for ideology estimation of each of these content subcategories. Finally, prior works do not undertake extensive performance benchmarking across multiple LLMs and diverse datasets. We utilize three different datasets (two containing news articles and one consisting of YouTube videos) and three different LLMs (open-source and closed-source) of various sizes. Our results shed light on which LLMs are better suited for ideology classification and what data/information augments LLM performance (e.g. whether additional video descriptions or channel information in prompts increases/reduces performance).
3 Problem Formulation
Our problem can be defined as follows: Given content that has an underlying ideological leaning (e.g., a YouTube video, a news article), we aim to determine if the content is liberal, neutral, or conservative. In this section, we describe our proposed approach where we utilize set coverage and balanced shot selection for ideology classification.
3.1 Proposed Approach
Algorithm 1 provides an overview of our approach. Given a LLM , a training set of content with ideology class labels (Liberal, Neutral, Conservative) , and number of demonstrations , our task is to classify a given query (an online post, video, or news article) as either liberal, neutral, or conservative. Our approach first builds a smaller candidate set of potential demonstrations and computes a relevance score for each demonstration in the candidate set. From the sorted candidate set, we select demonstrations equally from across all three ideology classes, i.e., only demonstrations can be selected from each ideology class. Thus, we skip over demonstrations with higher scores if demonstrations have already been selected from the class they belong to. The instruction prompts used can be found in Appendix A.
Test Instance:
LLM:
Prompt:
Training Set:
Number of Demonstrations:
Selected Demonstrations
Set() Candidate set using coverage
Label counts
3.2 Query Specifications
We specify the query to be classified as an online post, video, or news article that has a title, source, and a description. An LLM can leverage information from all three components to aid in classification, however, limitations such as context-window sizes and longer compute times for longer text make it worthwhile to explore the minimum amount of information that provides the best accuracy. As such, our experiments rely on four combinations: title alone, title and source, title and description, and title, source, and description.
| Sample Text |
| Trump tweets COVID-19 video so outrageous it was banned from Facebook and Twitter |
| Selected Demonstrations |
| Twitter removes Trump retweet that falsely claimed COVID-19 ‘has a cure’ |
| Reporter to sue Michigan Gov. Whitmer for allegedly not releasing COVID-19 nursing home data |
| AU Health masters COVID-19 therapy but loses $20 million |
| Facebook to label COVID-19 vaccine posts to combat misinformation, help people find where to get the shot |
| Trump Gives Delusional COVID-19 Press Conference Like A Broken Man |
| Florida Scientist Says She Was Fired For Not Manipulating COVID-19 Data |
| Wisconsin pharmacist arrested for sabotaging COVID-19 vaccines |
| Facebook openly conspiring with FBI to punish Trump supporters for wrongthink |
3.3 Demonstration Selection
We use the Set-BSR approach by Gupta et al. (2023) to select the demonstrations. It uses BERTScore-Recall (BSR) to measure how well a single example covers the salient aspects of a test input by leveraging contextual embeddings and recall-based token matching (see Table 2 for an example on how salient topics are covered). To optimize for groups of examples, it introduces a Set Coverage metric that ensures all salient aspects are collectively covered while minimizing redundancy. We select the top- ranked demonstrations such that there is an equal number of demonstrations selected from each ideology class. A comparison of random and balanced demonstration selection can be found in Appendix D.
4 Results and Evaluation
4.1 Experimental Setup
4.1.1 Datasets
For our evaluation, we rely on three distinct datasets, which have been used in prior work on ideology classification. While in our main experiments the candidate set sizes are set to high values, we show that similar results can be obtained by smaller candidate set sizes as well in Appendix C.
YouTube Slant Dataset. We sample 645,187 videos from the YouTube dataset by Haroon et al. (2023) which estimates the ideology of videos using the Twitter-based slant estimation approach Le et al. (2017); Barberá (2015); Wojcieszak et al. (2023) on a scale of -1 (most liberal) to +1 (most conservative) and was collected between 2021 and 2022. We do not explicitly check whether these videos or their corresponding channels are regional to the US but only that they are in English. A BERT-based political classifier Askari et al. (2024) is used to further identify 49,894 political videos to create the YouTube slant dataset, where we label videos with slant as liberal, as conservative, and videos in-between as neutral. The training set used for selection consists of 48,895 videos from which 20,000 candidates were selected and ranked using the Set-BSR approach. The test set consists of 1,000 videos sampled as to maintain an equal balance of ideologies. Overall, we have videos from over 8,523 channels across the entire training and testing sets (See Appendix G for where the slant cut-offs lie respective to popular news channels).
Ad Fontes News Articles. This dataset consists of news articles in the US that have been labeled by expert human annotators on Ad Fontes Media Bias Panel Otero (2019) from “most extreme left” to “most extreme right.” At least three experts rate each article from across the US political spectrum on a scale of -42 to +42 corresponding to extremely left and extremely right news articles, respectively. Additionally, Ad Fontes labels consider articles with scores lower than -14.0 to skew left and articles with scores greater than +14.0 as skewing right. Borrowing these cut-offs, we label articles with scores lower than -14.0 as Liberal, scores greater than +14.0 as Conservative, and everything in-between as Neutral. The training set consists of 33,396 news articles from which we build a candidate set of 8,000 using the Set-BSR approach. The test set contains 1,000 news articles with a balanced distribution of ideology.
AllSides News Articles. The dataset contains expert human annotations from AllSides444http://www.allsides.com/media-bias/, which relies on several human annotation approaches to rank news articles in the US such as blind surveys, expert reviews (on a scale of -6 to +6), and independent reviews. We use the training and test set of news articles used by Baly et al. (2019) from this dataset. The training set consists of 26,590 news articles, from which we build a candidate set of 8,000 using the Set-BSR approach. The test set consists of 1,300 videos.
4.1.2 LLMs
We use the following LLMs for our experiments:
OpenAI’s GPT-4o. We use the pre-trained GPT-4o model by OpenAI Hurst et al. (2024) using a paid API subscription.
Meta’s Llama-2. We use the HuggingFace library to run Meta’s Llama-2 model with 13B Touvron et al. (2023) parameters on a remote server with a RTX A6000 GPU, 256GB memory, and 48 cores.
MistralAI’s Mistral. We use the HuggingFace library to run MistralAI’s Mistral model with 7B parameters Jiang et al. (2023) on the same machine.
4.1.3 Baselines
We use two standard classification approaches as our supervised learning baselines.
Multilayer Perceptron (MLP). This model is a fully connected neural network that takes a 384-dimension text embedding as input to an initial dense hidden layer of 512 neurons with the tanh activation function. The final layer uses a softmax activation function over three outputs, providing a probability distribution over the three class labels. We use the Adam optimizer and categorical cross-entropy as the loss function.
RoBERTa. We use the base RoBERTa transformer model on the three training datasets by using its tokenizer to first embed the given text, and then training the model on the sequence classification task by using the ideology class labels. The model is fine-tuned for 50 epochs.
4.2 Results
Baselines. The baseline MLP model performed the worst, achieving an accuracy of , , and for the YouTube, Ad Fontes, and AllSides datasets when only the title was provided. The RoBERTa baseline achieved a score of on the YouTube Dataset and on the AllSides News dataset, marked improvements from the MLP model, but fared worse on the Ad Fontes dataset, with an accuracy of . This clearly shows that simpler models are not well-suited to the task of ideological estimation, especially when compared to more advanced transformer-based architectures. In Appendix B, we discuss different ablations of title, source, and description on the two baselines to see performance differences.
Zero-shot performance. Figure 1 shows the performance of the three LLMs and baselines across the three datasets. Focusing on the zero-shot setting, we notice that GPT-4o outperformed both Mistral and Llama2, achieving an accuracy of , , and across the YouTube, Ad Fontes, and AllSides datasets. Mistral performed better than Llama2, achieving a accuracy on the YouTube dataset closer to GPT-4o whereas Llama2’s accuracy was very low at . On the Ad Fontes dataset, GPT-4o performed better than Mistral at accuracy compared to Mistral’s . Finally, all LLMs performed poorly on the AllSides dataset with GPT-4o performing the best at , lower than the baseline RoBERTa model (fine-tuned on titles only). Comparing with the baselines, RoBERTa performed similar to GPT-4o at and on the YouTube Slant and AllSides datasets, respectively but far worse on the Ad Fontes dataset at . In contrast, the MLP performed the worst in the YouTube and Ad Fontes datasets, and on par with Mistral and Llama on the AllSides dataset.
Few-shot ICL comparison. Providing the set of demonstrations selected using the Set-BSR approach to the prompt , Figure 1 also shows how the accuracy of the models improved by increasing the number of demonstrations . The trend of GPT-4o outperforming Llama2 and Mistral persisted as the number of demonstrations increased. GPT-4o consistently achieved the highest scores across all datasets, suggesting that it is particularly well-suited for ideological estimation tasks. Mistral and GPT-4o showed a steady increase in performance as the number of demonstrations increased, both seeing a and increase in accuracy going from ( to for GPT-4o, and to for Mistral). Llama2 saw significant increases in accuracy across all datasets from , but grew worse as . Few-shot results for other dataset variations can be found in Appendix F.
Including Source and Description increases accuracy at times but not always. Augmenting the test instance to include source and description increased the performance of the models across all the datasets in the zero-shot setting. Figure 2 shows how model performance was affected by altering to test different combinations of the title, source, and description. We consider the best performing ablation of the baseline MLP and RoBERTa models as described in Appendix B. We observed that the source alone was a strong indicator of ideology in the two news datasets and less so in the YouTube slant dataset owing to the high accuracies. These two datasets comprise news articles from well-known news sources of whom the LLMs are most likely aware of. The model can then simply predict ideology based on its knowledge of the source’s ideology indicating a high correlation between the news articles and their sources considered. The accuracy is poorer in the YouTube slant dataset where the sources can be niche and less popular YouTube channels of whom the LLM may not be aware of and performed better when the title and description were provided. For the YouTube dataset, performance for GPT-4o increased by when the source and description was provided along with the title. Llama2 saw an increase of and Mistral saw an increase of when the title and source were provided. When augmenting the title with the source and description, we see significantly higher increases for Ad Fontes, with the performance of GPT-4o increasing by , Llama2 by , and Mistral by . Similar increases are seen for AllSides, where adding just the source pushed the accuracy of the models further. We present and discuss these ablations in more detail in Appendix B, and discuss edge cases where the ideology of the source can be misleading in Appendix E.
Non-political content is difficult to classify. Our findings so far have focused on exclusively political content. The improvement in the accuracy of the LLMs when the source is provided suggests that the source carries substantial weight in determining the ideology of content, where the title alone was previously misclassified. This improvement is more evident in the Ad Fontes and AllSides news datasets, where the sources are well-known news organizations, and less so in the YouTube slant dataset, where the sources can be any YouTube channel. To pinpoint the relative role of the title and source in determining the classification, we re-sample the entire YouTube slant dataset to procure three datasets: 1) political content from well-known news channels (political news), 2) non-political content from well-known news channels (non-political news), and 3) political news from non-news channels, such as influencers or pundits (political non-news). We utilize the BERT-based political classifier to identify political videos Askari et al. (2024) and use an extensive list of 1,625 news YouTube channels by Yu et al. (2024) to identify news channels. The same training as the YouTube slant dataset is used and ideology is estimated using GPT-4o.
| Political | Political | Non-political | |
| news | non-news | news | |
| Liberal | 314 | 309 | 125 |
| Neutral | 435 | 436 | 207 |
| Conservative | 251 | 255 | 76 |
| Total | 1,000 | 1,000 | 408 |
Breaking this down by the class labels, Figure 3 shows the change in prediction accuracy after adding source information using a heatmap of changes across the three datasets. The heatmap shows which corresponding class was a given label misclassified as before providing the source allowing us to analyze the influence of the source when estimating ideology. Coupling this with qualitative examples from Appendix H provides insight into the influence of the source in classifying ideology.
For political news, we see an improvement in accuracy of content previously misclassified as neutral to liberal (+) and conservative (+), becoming almost more accurate in both cases. When looking at the kinds of titles re-classified in Table 13 in Appendix H, we see that the titles themselves do not indicate the any ideological leaning. For example, “Haberman says Trump has been diminished but isn’t dead politically” was classified as neutral but when the LLM was informed it was from CNN, it was re-classified as liberal. Similarly, the video titled “Transgender Clinical Psychologist on people de-transitioning | Dr. Erica Anderson” was re-classified as conservative when the LLM learned it came from Newsmax, a conservative outlet, despite the title itself indicating no such political leaning. Cases where the label changed from liberal to conservative or vice-versa showed more nuance. The video titled “Democrats introduce bill to block Trump from holding office” was classified as liberal but then re-classified as conservative when the source Newsmax was provided which was consistent with the content of the video. This is true for the non-political news dataset as well, where the titles do not depict any political inclination as can be seen in tab:qual-nonpolnews in Appendix H but the model re-classified it according to the ideology of the news source. Finally, for the political non-news videos, we do not see much movement as the LLM is likely unfamiliar with the sources provided. The cases where we do see movement as seen in Table 15, it is also driven by the source of the video as well. We see some movement towards the neutral as well. For example, the videos titled “The Trouble with the Electoral College” and “Is capitalism actually broken?” are liberal in ideology but when the model learns they are from educative channels CGP Grey and TED-Ed, it re-classifies them as neutral.
4.3 Discussion
Leveraging few-shot ICL and additional metadata for best performance. Combining the few-shot ICL and incorporating the source resulted in the highest gains in GPT-4o’s performance across all three datasets. The model achieved a increase in the YouTube slant dataset (), 18% increase in the Ad Fontes dataset (), and 37% increase in the AllSides dataset (). These improvements demonstrate how GPT-4o leverages details from relevant demonstrations and its existing knowledge of news sources to more accurately reason about the ideology of content.
Multi-modal classification using YouTube thumbnails does not improve GPT-4o’s performance. We experiment with adding video thumbnails as part of the prompt to GPT-4o. Despite the potential of multi-modal classification methods to enhance model performance by incorporating visual information, our experiments reveal that using YouTube thumbnails as an additional modality does not significantly improve GPT-4o’s performance. When comparing the results of GPT-4o with and without the inclusion of YouTube thumbnails, the accuracy of the model dropped to from on the YouTube slant dataset. This result indicates that, for the problem of ideology estimation from YouTube content, visual information does not contribute additional meaningful signals beyond what is captured in the text and instead can be misleading for the model. However, more extensive analysis and experiments are needed for incorporating multi-modal information for ideology classification and can be undertaken in future work.
Chain-of-thought reasoning is ineffective. We apply chain-of-thought (CoT) reasoning to GPT-4o for our task by prompting the model to “try to think step-by-step to reason about the ideology of the post”. We observed that this did not significantly improve accuracy despite its known benefits Kojima et al. (2024). On the YouTube slant dataset, the model accuracy dropped from to when CoT reasoning was used. One explanation for this finding is that ideology estimation is primarily reliant on identifying patterns in content that are explicit rather than requiring deep logical steps. The task might not demand the same kind of multi-step reasoning that CoT typically helps with in applications, such as arithmetic or procedural problem-solving.
Accuracy of LLMs is comparable with human annotation. The high accuracy of GPT-4o is comparable to human annotation on the AllSides and Ad Fontes News datasets, both of which are human-annotated. On these, GPT-4o achieved high accuracy, closely aligning with human judgments, exceeding in some configurations. For additional validation, a domain expert annotated a sample of 561 videos in the YouTube dataset, finding that GPT-4o agreed with the human annotation of the time. This suggests that GPT-4o is able to effectively mirror human-level ideological classification, even in more informal content. The model’s performance highlights its capacity to generalize patterns and signals used by human annotators, making it a valuable tool for automated ideological estimation.
5 Conclusion
This study explored the efficacy of few-shot learning using Large Language Models for the task of ideology estimation across diverse textual datasets, including news articles and YouTube videos. Our findings demonstrate that LLMs, particularly GPT-4o, achieve state-of-the-art performance, surpassing traditional supervised baselines. By leveraging advanced few-shot in-context learning methods, we addressed challenges of scalability and robustness, showcasing the ability of LLMs to adapt to evolving contexts with minimal supervision.
6 Limitations
Despite the promising results and contributions outlined in this paper, several limitations must be acknowledged to identify areas for future work. The performance of our approach is heavily reliant on the capabilities of pre-trained LLMs such as GPT-4o, Llama2-13B, and Mistral-7B. Variability in model updates or differences in underlying pre-training data could affect reproducibility and consistency across different deployments. In addition, while incorporating metadata, such as source information improved performance, it also introduced the potential for biased predictions. Models may disproportionately rely on known ideological leanings of sources, potentially misclassifying nuanced or neutral content.
Furthermore, our experiments focus on text, and extending these methods to classify ideology in non-textual data such as video or audio remains an open challenge and an important direction for future work. Lastly, we acknowledge that the datasets used for evaluation, while diverse, may not fully capture the breadth of content and ideological nuance present in real-world social media platforms. For instance, the balance of liberal, neutral, and conservative content in the datasets may not reflect natural distributions, and also the majority of content consumed by users is generated by other users themselves, not by news media organizations or YouTube channels, suggesting the need to adapt our approach to user-generated posts and comments.
7 Ethics Statement
Automated ideology estimation carries inherent risks, including misuse for surveillance or censorship. Furthermore, reliance on LLMs can perpetuate biases present in training data, which might influence predictions in unintended ways. While the study evaluates multiple datasets, the generalizability of the approach across all social media platforms—with varying content types and user behaviors–—is yet to be fully established.
Addressing these limitations could involve refining LLM prompting techniques, improving model transparency, exploring alternative datasets, and incorporating mechanisms for bias mitigation and real-time adaptability. Future research should also emphasize ethical frameworks for deploying such systems responsibly, ensuring that advancements in technology contribute positively to societal discourse without exacerbating existing challenges.
Appendix
Appendix A Prompts used for the LLM
Here, we present the prompts that are used for querying the models.
Prompt for querying with just the title.
Classify the following news article titles as ideologically liberal, neutral, or conservative. Titles with no ideological content are classified as neutral. Only respond with the final answer.
Prompt for querying with the title and source.
Classify the following news article titles as ideologically liberal, neutral, or conservative. Titles with no ideological content are classified as neutral. The news source is also specified for additional context. Only respond with the final answer.
Prompt for querying with the title, description, and source.
Classify the following news article titles as ideologically liberal, neutral, or conservative. Titles with no ideological content are classified as neutral. The news description is also specified for additional context. Only respond with the final answer.
Appendix B Ablations on different parameters
In this section, we report on the various ablations of title, source, description, and number of shots for the various baseline and LLM models.
B.1 MLP Baseline
Table 4 shows the accuracy of the MLP model under different title, source, and description ablations. The multi-layer perceptron model achieved good accuracy when the source was provided. In the AllSides news dataset, wherein we observe that the source is highly correlated with article ideology, the model achieves accuracy when just the source is provided. Similarly, MLP achieves accuracy on the YouTube slant dataset when title, source, and article description are provided. Finally, the model achieves a high accuracy of on the Ad Fontes news dataset when just the title and source are provided.
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Desc. | Slant | News | News |
| ✓ | 0.51 | 0.38 | 0.43 | ||
| ✓ | 0.61 | 0.55 | 0.69 | ||
| ✓ | 0.55 | 0.38 | 0.47 | ||
| ✓ | ✓ | 0.63 | 0.62 | 0.57 | |
| ✓ | ✓ | 0.57 | 0.43 | 0.46 | |
| ✓ | ✓ | 0.62 | 0.60 | 0.58 | |
| ✓ | ✓ | ✓ | 0.64 | 0.61 | 0.52 |
B.2 RoBERTa baseline
Table 5 shows the accuracy of the RoBERTa model under different title, source, and description ablations. The RoBERTa model performed well on the YouTube slant dataset achieving an accuracy of when title, source, and description were all provided. The performance was worse on the news datasets, with a maximum of on the Ad Fontes news dataset and on the AllSides news dataset.
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Desc. | Slant | News | News |
| ✓ | 0.57 | 0.35 | 0.52 | ||
| ✓ | 0.45 | 0.48 | 0.48 | ||
| ✓ | 0.62 | 0.35 | 0.50 | ||
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | 0.64 | 0.38 | 0.53 | |
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | ✓ | 0.65 | 0.44 | 0.55 |
B.3 GPT-4o
Table 6 shows the accuracy of the GPT-4o model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best when given a higher number of demonstrations for the YouTube slant and Ad Fontes news datasets. For the YouTube dataset, the model performed best when title, source, and description were all provided whereas for the Ad Fontes news dataset, only title and source were sufficient. The AllSides news dataset showed a different trend in that zero demonstrations gave the best score when only the source was provided. We observed in the test set for this dataset that source and target label were highly correlated (i.e., a video from Fox, a well-known conservative outlet, was almost always labeled conservative) making it simpler for the model to just assign the source ideology as the target ideology. In Appendix E, we investigate this further by considering edge cases in the test set.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.64 | 0.69 | 0.50 | ||
| 0 | ✓ | 0.63 | 0.80 | 0.88 | ||
| 0 | ✓ | 0.56 | 0.64 | 0.44 | ||
| 0 | ✓ | ✓ | 0.64 | 0.78 | 0.74 | |
| 0 | ✓ | ✓ | 0.63 | 0.75 | 0.55 | |
| 0 | ✓ | ✓ | 0.59 | 0.82 | 0.85 | |
| 0 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.72 |
| 4 | ✓ | 0.67 | 0.74 | 0.55 | ||
| 4 | ✓ | ✓ | 0.73 | 0.87 | 0.83 | |
| 4 | ✓ | ✓ | 0.69 | 0.80 | 0.58 | |
| 4 | ✓ | ✓ | ✓ | 0.72 | 0.86 | 0.82 |
| 8 | ✓ | 0.68 | 0.73 | 0.55 | ||
| 8 | ✓ | ✓ | 0.72 | 0.87 | 0.86 | |
| 8 | ✓ | ✓ | 0.70 | 0.81 | 0.59 | |
| 8 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.84 |
| 12 | ✓ | 0.69 | 0.74 | 0.55 | ||
| 12 | ✓ | ✓ | 0.74 | 0.87 | 0.87 | |
| 12 | ✓ | ✓ | 0.71 | 0.82 | 0.59 | |
| 12 | ✓ | ✓ | ✓ | 0.71 | 0.87 | 0.84 |
B.4 Llama2
Table 7 shows the accuracy of the Llama2 model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best at 4 or fewer demonstrations indicating that more demonstrations leads to a drop in performance likely due to a limited context window. For the YouTube and Ad Fontes news datasets, the model performed best when title, source, and only four demonstrations were provided. The AllSides news dataset showed similar trends as GPT-4o in that the source was sufficient without any demonstrations to achieve the best accuracy.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.53 | 0.49 | 0.42 | ||
| 0 | ✓ | 0.61 | 0.52 | 0.82 | ||
| 0 | ✓ | 0.52 | 0.49 | 0.37 | ||
| 0 | ✓ | ✓ | 0.63 | 0.57 | 0.65 | |
| 0 | ✓ | ✓ | 0.58 | 0.57 | 0.46 | |
| 0 | ✓ | ✓ | 0.48 | 0.65 | 0.71 | |
| 0 | ✓ | ✓ | ✓ | 0.58 | 0.69 | 0.59 |
| 4 | ✓ | 0.64 | 0.60 | 0.54 | ||
| 4 | ✓ | ✓ | 0.68 | 0.79 | 0.75 | |
| 4 | ✓ | ✓ | 0.66 | 0.68 | 0.51 | |
| 4 | ✓ | ✓ | ✓ | 0.64 | 0.79 | 0.68 |
| 8 | ✓ | 0.60 | 0.56 | 0.51 | ||
| 8 | ✓ | ✓ | 0.63 | 0.76 | 0.74 | |
| 8 | ✓ | ✓ | 0.63 | 0.67 | 0.52 | |
| 8 | ✓ | ✓ | ✓ | 0.60 | 0.78 | 0.67 |
| 12 | ✓ | 0.60 | 0.60 | 0.50 | ||
| 12 | ✓ | ✓ | 0.62 | 0.75 | 0.72 | |
| 12 | ✓ | ✓ | 0.58 | 0.68 | 0.49 | |
| 12 | ✓ | ✓ | ✓ | 0.56 | 0.78 | 0.67 |
B.5 Mistral
Table 8 shows the accuracy of the Mistral model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best at zero demonstrations indicating that more demonstrations leads to a drop in performance likely due to a limited context window. For the YouTube and Ad Fontes news datasets, the model performed best when title, source, and description were provided with zero demonstrations. The AllSides news dataset again showed similar trends as GPT-4o in that the source was sufficient without any demonstrations to achieve the best accuracy.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.61 | 0.57 | 0.44 | ||
| 0 | ✓ | 0.61 | 0.72 | 0.83 | ||
| 0 | ✓ | 0.52 | 0.59 | 0.34 | ||
| 0 | ✓ | ✓ | 0.67 | 0.72 | 0.71 | |
| 0 | ✓ | ✓ | 0.62 | 0.62 | 0.41 | |
| 0 | ✓ | ✓ | 0.57 | 0.73 | 0.66 | |
| 0 | ✓ | ✓ | ✓ | 0.65 | 0.74 | 0.58 |
| 4 | ✓ | 0.61 | 0.57 | 0.42 | ||
| 4 | ✓ | ✓ | 0.65 | 0.73 | 0.73 | |
| 4 | ✓ | ✓ | 0.62 | 0.61 | 0.44 | |
| 4 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.60 |
| 8 | ✓ | 0.62 | 0.59 | 0.44 | ||
| 8 | ✓ | ✓ | 0.64 | 0.71 | 0.73 | |
| 8 | ✓ | ✓ | 0.63 | 0.62 | 0.48 | |
| 8 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.61 |
| 12 | ✓ | 0.64 | 0.61 | 0.44 | ||
| 12 | ✓ | ✓ | 0.66 | 0.71 | 0.73 | |
| 12 | ✓ | ✓ | 0.61 | 0.63 | 0.48 | |
| 12 | ✓ | ✓ | ✓ | 0.63 | 0.69 | 0.61 |
Appendix C Candidate set size can be smaller
One might argue that the candidate set from which demonstrations are selected is similar to the training set used for training supervised models and, thus, a large amount of annotated data is also required for LLMs. While the size of the candidate set was configured high across all three datasets, we further analyze whether the need for annotated data can be alleviated by considering a smaller candidate set size as well. Specifically, we set the candidate set size to only 500 samples, a fraction of what is generally required to train supervised models. Testing this across all three LLMs and all three datasets in Table 9, we observe that the accuracy is not majorly effected by this change. This is likely because the Set-BSR approach already selects the most optimal demonstrations that maximize coverage.
| LLM | YouTube Slant | Ad Fontes News | AllSides News |
| GPT-4o | 0.66 | 0.74 | 0.52 |
| Llama2 | 0.60 | 0.57 | 0.50 |
| Mistral | 0.63 | 0.59 | 0.42 |
Appendix D Balanced demonstration selection is better than random
We experiment with two demonstration selection techniques namely 1) Random where the demonstrations are randomly selected from the candidate set and 2) Balanced BERTScore where the BERTScore-Recall metric is used to select demonstrations in a label-balanced manner (equal liberal, neutral, and conservative demonstrations). Comparing the performance of Mistral with 8 demonstrations, we see in Table 10 that there is an average increase of in accuracy across the datasets when using Balanced BERTScore selection over random.
| Dataset | Random | Balanced BERTScore |
| YouTube Slant | 0.52 | 0.61 |
| Ad Fontes News | 0.49 | 0.57 |
| AllSides News | 0.39 | 0.42 |
Appendix E Sources can be misleading
Our prior results showed that providing the source information significantly increases the accuracy. However, not all content from a given source may share the ideology of the source itself Haroon et al. (2023), as some liberal articles can come from conservative outlets. To highlight this particular scenario, we identify cases in our test set where the ideology of the article or video is different from the ideology of its source for all three datasets (e.g., a liberal article from Fox News or a conservative article from MSNBC). We rely on AllSides’ media bias chart555https://www.allsides.com/media-bias/media-bias-chart for source ideology and map these sources to their corresponding news articles and YouTube videos. We then compare the accuracy of GPT-4o when 1) only title is provided, 2) only source is provided, and 3) both title and source are provided.
| Liberal Sources | Conservative Sources |
| ABC News | Breitbart |
| Associated Press | CBN |
| CBC News | Daily Caller |
| CBS News | Daily Mail |
| CNN | Daily Wire |
| Daily Beast | Fox Business |
| HuffPost | Fox News |
| Jacobin | New York Post |
| MSNBC | Newsmax |
| Mother Jones | The American Conservative |
| NPR | The American Spectator |
| Slate | The Blaze |
| The Atlantic | The Daily Caller |
| The Guardian | The Epoch Times |
| The Intercept | The Federalist |
| The Nation | The Federalist Society |
| The New York Times | The Post Millennial |
| The New Yorker | The Washington Free Beacon |
| Vox | Washington Examiner |
Table 11 lists the names of the left and right sources identified and mapped to the three datasets. In the YouTube test set, we identified 40 cases where an neutral/conservative article from a liberal source and 21 cases where a liberal/neutral article came from a conservative source. The Ad Fontes news test set had 38 cases of neutral/conservative articles from liberal sources and 11 cases of liberal/neutral articles from conservative sources. Finally, the AllSides news test set had 98 neutral/conservative articles from liberal sources and 0 liberal/neutral articles from conservative sources.
| Non-liberal articles from liberal sources | ||||
| Title | Source | YouTube Slant | Ad Fontes News | AllSides News |
| ✓ | 0.63 | 0.58 | 0.49 | |
| ✓ | 0.47 | 0.38 | 1.00 | |
| ✓ | ✓ | 0.66 | 0.70 | 1.00 |
| Non-conservative articles from conservative sources | ||||
| Title | Source | YouTube Slant | Ad Fontes News | AllSides News |
| ✓ | 0.45 | 0.57 | — | |
| ✓ | 0.00 | 0.00 | — | |
| ✓ | ✓ | 0.00 | 0.52 | — |
Table 12 reports the accuracy of the LLMs on these specific test cases for liberal and conservative sources respectively. We see that the LLMs suffer from incorrect classifications when only the source is considered as the ground truth contradicts the ideology of the source. Here, providing the title increases the accuracy significantly especially in the conservative sources where the accuracy on just the source is for the YouTube and Ad Fontes datasets. In fact, providing the title for the YouTube test set in this case still does not improve the accuracy seeing as the model accuracy is still . This suggests that the LLM overtly relies on the source in this case completely disregarding title wherein accuracy on just the title was close to . Similar trends can be seen for the Ad Fontes news test set where the accuracy of source alone was and title was , and both was . We observe that the AllSides news test set still has better performance when the source is provided at a . On closer inspection, we observe that in the test set sampled, there is a high correlation between the source and article ideology which likely leads to such a high performance when source is provided.
Appendix F Few-shot results for additional datasets
Figures 4, 5, and 6 show the increases in accuracy across all datasets by increasing the number of shots.
Appendix G Placement of popular news channels based on the slant cut-offs
Figure 7 shows the placement of popular news channels based on the YouTube slant dataset. Corresponding cut-offs at -0.33 and +0.33 are also shown for the liberal and conservative respectively.
Appendix H Qualitative results for the political news analysis for YouTube videos
Qualitative results for political news videos, non-political news videos, and political non-news videos that changed predictions as a result of adding information about the source are provided in Tables 13, 14, and 15 respectively.
| Title | Channel |
| Liberal to Neutral | |
| How welcome are Russian deserters in Georgia? | Focus on Europe | DW News |
| Ukrainians ditch Russian traditions and celebrate Christmas with the West | Sky News |
| Asian Americans Debate Model Minority & Asian Hate | VICE Debates | VICE |
| Secret Service under scrutiny ahead January 6 committee’s final report | PBS NewsHour |
| White House Correspondent Jeff Mason on his tense exchanges with Trump | Reuters |
| Liberal to Conservative | |
| Democrats introduce bill to block Trump from holding office | Newsmax |
| LIVE: Jan. 6 Committee Votes on Criminal Referrals, Final Report | NTD |
| New CRIMINAL REFERRALS Against Former President Trump Unveiled by Jan. 6th Committee | BlazeTV |
| Tonight’s Tipping Points: Hate Crimes, Planned Parenthood, & Trans Military Members! | One America News Network |
| Trolling #AMFEST2022 Attendees as Left-Winger ’Garth Griffin’ of Edge News | Project Veritas |
| Neutral to Liberal | |
| Haberman says Trump has been diminished but isn’t dead politically | CNN |
| Kevin McCarthy’s Path To Speaker Is In Peril | MSNBC |
| Anger grows in Russia over Ukraine’s New Year’s Day strike on Russian troops | CNN |
| Rep. Adam Kinzinger Delivers Final Speech As Congressman | MSNBC |
| ’Shocked’: Russia expert explains why Putin’s change of word is significant | CNN |
| Neutral to Conservative | |
| UK Weather: What’s behind Britain’s big freeze? | UK snow chaos explained | Daily Mail |
| Transgender Clinical Psychologist on people de-transitioning | Dr. Erica Anderson | Newsmax |
| Actor slams crypto as ’biggest Ponzi scheme ever’ during Senate hearing | Fox News |
| Rep. Spartz explains her reasoning for voting ’present’ for Speaker | Fox News |
| Internet users allege Bryan Kohberger attended Idaho victims’ vigil | Fox News |
| Conservative to Liberal | |
| Trump Supporter: Too Much Diversity Will Kill America | The Young Turks |
| Russian foreign minister: Ukraine must give up occupied territories | CNN |
| These Trump Supporters Are CONVINCED "Sleepy Joe Poopy Pants Isn’t The Acting President" | The Young Turks |
| Steve Bannon reacts to Trump’s ’major announcement’ | CNN |
| Trump adviser tells reporter ’the magic is gone’ in new article about re-election bid | CNN |
| Conservative to Neutral | |
| Justin Amash APPLAUDS Revolt Against McCarthy, Says Ruling Oligarchy CRUSHES Dissent | The Hill |
| State Of Emergency Declared At Texas Border | NBC News |
| Watch Trump claim the army retook airports during the Revolutionary War in July Fourth speech | The Globe and Mail |
| Trump Calls Jan. 6 Committee A ’Kangaroo Court’ After Criminal Referrals | NBC News |
| ’They shot up the WRONG house!’ Sheriff Grady Judd on young kids shot at while in bed in Polk County | WFLA News Channel 8 |
| Title | Channel |
| Liberal to Neutral | |
| TRIPLEDEMIC Or Free Market Failure To Blame For Medicine SHORTAGE? STRUGGLE To Find Children’s Meds | The Hill |
| Watch: Jan. 6 Committee Presents Key Video Evidence Compilation | NBC News |
| Neutral to Liberal | |
| Top U.S. & World Headlines — December 27, 2022 | Democracy Now! |
| Top U.S. & World Headlines — January 3, 2023 | Democracy Now! |
| Raskin On Cancer Diagnosis: ’Totally My Plan To Make It Through This Thing’ | MSNBC |
| Watch The 11th Hour Highlights: Dec. 20 | MSNBC |
| Plans For Zelenskyy D.C. Trip Came Together In A ’Matter Of Days’ | MSNBC |
| Neutral to Conservative | |
| McCarthy speakership battle: What happens if he concedes? | Fox Business |
| The yule log Christmas tradition | Fox News |
| “Woman” Is Named Word of the Year | DailyWire+ |
| What is the controversy surrounding Meghan Markle and her sister? | Fox News |
| ’DOUBLE WHAMMY’: Real estate expert issues pricing warning | Fox Business |
| Conservative to Liberal | |
| All Gestures, All Shock Opera: DeSantis Launches His Newest Battle | MSNBC |
| Christian Influencer Makes Surprising Confession About Trans Porn | The Young Turks |
| The War on Christmas RAGES ON! | The Young Turks |
| Alex Jones Says He’s Paying His Personal Trainer $100,000 A Week | The Young Turks |
| ’Look at where he is today’: Bash on Kinzinger’s farewell speech | CNN |
| Conservative to Neutral | |
| Cancelling the F-22 Fighter is Worse than You Think | Task & Purpose |
| Title | Channel |
| Liberal to Neutral | |
| Is capitalism actually broken? | TED-Ed |
| Update from Ukraine | Ukraine is getting ready for the winter Attack | Ruzzia will lose it | Denys Davydov |
| Van Lathan: A Jewish Producer Confronted Kanye at TMZ when Ye Said He Loves Hitler (Part 2) | djvlad |
| The Trouble with the Electoral College | CGP Grey |
| Those Times Sam Harris Made Us Slap Our Knees | Enlightainment |
| Liberal to Conservative | |
| The New Beginning of Same-Sex Marriage | Ben Shapiro |
| Meet Kamala Harris’s 6-Year-Old Speechwriter | The Babylon Bee |
| Pro-Choice Bosses - A Generous Offer To Our Pregnant Employees | Live Action |
| Confronting Charlie Kirk and Tim Pool at Turning Point USA America Fest | Alex Stein |
| I Showed Up To An Anti-Gun Protest In Los Angeles | Let’s Talk | James Klug |
| Neutral to Liberal | |
| Meanwhile… Pantone’s Color of the Year | Macron Makes Condoms Free In France | The Late Show with Stephen Colbert |
| Grandma Arrested For Feeding The Homeless?! Danbury Court Case Clarifcation + GIVEAWAY! | Long Island Audit |
| LIVE: EMERGENCY January 6 Committee hearing - Day 6 | Brian Tyler Cohen |
| The Funniest Donald Trump Stories On The Graham Norton Show | Part One | The Graham Norton Show |
| Prison Officer BUSTED After Massive Nazi Gang Attack Caught | The Damage Report |
| Neutral to Conservative | |
| Winning With The Hand You’re Dealt with Rick Warren | Saddleback Church |
| Fighting the Darkest Criminal Enterprise on the Planet | The 700 Club |
| STORE WARS: Target, Walmart & Home Depot go to war against Organized Retail theft | News For Reasonable People |
| Warner MAY BE SUED over Amber Heard and Aquaman 2?! NEW Hollywood Lawsuit! | ThatUmbrellaGuy |
| Historical alteration on the Ukraine Map: Russian 1st Corps withdrawing en masse from 4 locations! | Divine Justice |
| Conservative to Liberal | |
| Elon Musk Finally CRACKS In Ultimate Outburst Over Free Speech | The Damage Report |
| Chris Wallace: This Would Never Have Happened Under Nancy Pelosi | The Late Show with Stephen Colbert |
| Mike Lindell Snaps After Hannity Admits There Was No Fraud | Tommy Campbell |
| Trump ATTACKS the FBI and DOJ and Loses his Mind in NEW POSTS | MeidasTouch |
| BREAKING: Kari Lake slammed with brutal legal news | Brian Tyler Cohen |
| Conservative to Neutral | |
| Why US Let Russia Become the Biggest Nuclear Power | The Infographics Show |
| The CIA’s Biggest Scam Exposed | KiraTV |
| Cop Pulls Over "Legal Expert" And SCHOOLS HIM On The Law | Audit the Audit |
| Why US Enemies are Scared of Ghosts (Elite Special Forces) | The Infographics Show |
| "INSPIRATIONAL HOLIDAY VIDEO" — A Bad Lip Reading of Joe Biden | Bad Lip Reading |
Appendix I Code and Reproducibility
Our code and data used for prompting the models and analyzing the results can be found here: https://anonymous.4open.science/r/llm-ideology-FC41/