### v2

# “Whose Side Are You On?” Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection

Abstract
Existing approaches to classifying ideology of online content are limited in that they require extensive human effort, labeling of large datasets, and are not able to adapt to evolving ideological contexts. This paper explores the potential of Large Language Models (LLMs) for classifying the political ideology of online content in the context of the two-party US political spectrum through in-context learning (ICL). Our extensive experiments involving demonstration selection in label-balanced fashion, conducted on three datasets comprising news articles and YouTube videos, reveal that our approach significantly outperforms zero-shot and traditional supervised methods. Additionally, we evaluate the influence of metadata (e.g., content source and descriptions) on ideological classification and discuss its implications. Finally, we show how providing the source and description influences the LLM’s classification.
“Whose Side Are You On?”
Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection
Muhammad Haroon University of California, Davis mharoon@ucdavis.edu Magdalena Wojcieszak University of California, Davis mwojcieszak@ucdavis.edu Anshuman Chhabra University of South Florida anshumanc@usf.edu
1 Introduction
Social media platforms are one of the leading sources of information for the majority of Americans Barthel et al. (2020). Observers worry that content on these platforms can cause political extremism Tufekci (2018); Roose (2019), and facilitate the formation of ideological filter bubbles that restrict exposure to diverse viewpoints Pariser (2011). Empirical evidence for these worries is inconsistent Haroon et al. (2023); Hosseinmardi et al. (2021); Chen et al. (2023); Nyhan et al. (2023) in part because scholars use different methodologies (audits versus observational data) and in part because of a lack of robust (i.e., highly accurate across domains) and scalable (i.e., requiring minimal human supervision) methods for estimating political ideology of the content users consume.
In this paper, we advance on literature showing that Large Language Models (LLMs) can be applied to ideology estimation of text in the context of the two-party US political spectrum (democrat/liberal and republican/conservative). They can also adapt to new and evolving information by simple adjustments of the provided prompt. Thus, LLMs can serve as a compelling choice for robust and scalable ideology estimation across diverse text domains and platforms.
Past applications of LLMs in ideology estimation Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023) have primarily focused on zero-shot approaches, utilizing pre-trained LLMs without employing more nuanced few-shot in-context learning (ICL) strategies. Our work advances these prior studies by proposing a modified version of coverage-based ICL that selects demonstrations in a class-balanced manner, significantly improving performance over simple zero-shot prompting and coverage-based ICL. We expand on this by conducting an extensive study through rigorous experiments on three distinct datasets on the capabilities of various LLMs (Llama2-13B, Mistral-7B, and GPT-4o) in estimating the ideology of text. Furthermore, we analyze the effect additional metadata such as content source and description has on the classifications of the LLM. Our analyses provides critical insights into the limitations and potentials of LLMs for nuanced ideological analysis.
2 Related Work
Prior approaches to ideology estimation have generally fallen under three types. (1) Manual annotations by experts Otero (2019) which, although accurate and robust, are not scalable due to the ever-increasing volume of social media content. (2) Training supervised classifiers Baly et al. (2020) which, although highly scalable, is limited by a need for large volumes of human-labeled data and have difficulty adapting to dynamic and ever evolving ideological contexts (e.g., cross-platform estimation, emerging news events, changes in the political environment). (3) Utilizing social media which uses data from external platforms such as Twitter Le et al. (2017); Barberá (2015); Wojcieszak et al. (2023); Eady et al. (2020) or Reddit Lai et al. (2024) to assess how a given piece of content is shared among users of different ideologies to estimate the ideology of the content itself. This approach is limited in that it requires the content to have been shared by a representative number of users and is agnostic of the nature of the content itself. These limitations prevent these prior approaches from being robust and scalable. On the other hand, in contrast with prior work that uses LLMs for ideology classification mostly in a zero-shot manner Heseltine et al. (2024); Törnberg (2023); Gilardi et al. (2023), we introduce a class-balanced, coverage-based ICL framework that achieves significant performance improvements over the zero-shot setting.
3 Proposed Approach
Our research problem can be defined as follows: Given content that has an underlying ideological leaning (e.g., a YouTube video, a news article), we aim to determine if the content is liberal, neutral, or conservative. Next, we describe our proposed approach for LLM based ideology classification.
Algorithm 1 presents our proposed ICL solution framework. Given a LLM , a training set of content with ideology class labels (Liberal, Neutral, Conservative) , and number of demonstrations , our task is to classify a given query (an online post, video, or news article) as either liberal, neutral, or conservative. Our approach first builds a smaller candidate set of potential demonstrations and computes a relevance score for each demonstration in the candidate set. From the sorted candidate set, we select demonstrations equally from across all three ideology classes, i.e., only demonstrations can be selected from each ideology class. Thus, we skip over demonstrations with higher scores if demonstrations have already been selected from the class they belong to. The instruction prompts used can be found in Appendix A.
3.1 Test Query Specifications
We specify the test query to be classified as an online post, video, or news article that has a title, source, and a description. An LLM can leverage information from all three components to aid in classification, however, limitations such as context-window sizes and longer compute times for longer text make it worthwhile to explore the minimum amount of information that provides the best accuracy. As such, our experiments rely on four combinations: title; title and source; title and description; and title, source, and description.
3.2 Demonstration Selection
We select demonstrations for in-context learning (ICL) to augment LLM task performance. We utilize Set-Cover Vazirani (2003) to maximize coverage across the candidate pool, where BertScore-Recall (BSR) is chosen as the coverage metric. This strategy is known as Set-BSR, and has been shown to be successful for ICL with multiple classes (Gupta et al., 2023). We also ensure there is an equal number of demonstrations selected from each ideology. To validate this design choice, a comparison of random and balanced demonstration selection can be found in Appendix D.
Test Instance:
LLM:
Prompt:
Training Set:
Number of Demonstrations:
Selected Demonstrations
Set() Candidate set using coverage
Label counts
4 Experimental Setup
4.1 Datasets
For our evaluation, we rely on three distinct datasets, which have been used in prior work on ideology classification.
YouTube Slant. We use the 49,894 political videos from the YouTube dataset by Haroon et al. (2023) which estimates the ideology of videos using a Twitter-based slant estimation approach Barberá (2015) on a scale of -1 (most liberal) to +1 (most conservative). We label videos with slant less than -0.33 as Liberal, greater than +0.33 as Conservative, and in-between as Neutral. For the ICL demonstration pool, we uniformly randomly sample 20,000 candidates from the full test set. The test set consists of 1,000 videos sampled as to maintain an equal balance across ideologies in testing.
Ad Fontes News. This dataset consists of US news articles that have been labeled by expert human annotators on Ad Fontes Media Bias Panel Otero (2019). Articles with scores lower than -14.0 are labeled as Liberal, greater than +14.0 as Conservative, and in-between as Neutral. The training set consists of 33,396 news articles from which we build a candidate set of 8,000 using the Set-BSR approach and the test set contains 1,000 news articles with a balanced distribution of ideology.
AllSides News. The dataset contains expert human annotations from AllSides111http://www.allsides.com/media-bias/, which relies on several human annotation approaches to rank news articles in the US such as blind surveys, expert and independent reviews. We use the same training and test sets used by Baly et al. (2019). The training set consists of 26,590 news articles, from which we build a candidate set of 8,000 using the Set-BSR approach. The test set consists of 1,300 videos.
While in our main experiments the candidate set size is large, we show in Appendix C that smaller candidate set sizes yield similar results.
4.2 LLMs
We use the following LLMs for our experiments:
OpenAI GPT-4o. We use the pre-trained GPT-4o model by OpenAI Hurst et al. (2024).
Meta Llama-2-13B. We use HuggingFace to run Meta’s Llama-2 model with 13B Touvron et al. (2023) parameters on a remote server with a RTX A6000 GPU, 256GB memory, and 48 cores.
MistralAI Mistral-7B. We use HuggingFace to run MistralAI’s Mistral model with 7B parameters Jiang et al. (2023) on the same machine.
4.3 Baselines
We use two standard classification approaches as our supervised learning baselines.
Multilayer Perceptron (MLP). A fully connected neural network that takes a 384-dimension text embedding as input, a dense hidden layer of 512 neurons with the tanh activation function, and a three output softmax final layer (for three ideologies).
RoBERTa Baly et al. (2020). As in Baly et al. (2020), we use the base RoBERTa transformer model and fine-tune on each of the three full training datasets for 50 epochs. Unlike the LLM which only receives a small number of demonstrations, the RoBERTa model sees the full dataset(s).
5 Results
Baselines. The baseline MLP model performed the worst, achieving an accuracy of , , and for the YouTube, Ad Fontes, and AllSides datasets when only the title was provided. The RoBERTa baseline achieved a score of on the YouTube dataset and on the AllSides News dataset, marked improvements from the MLP model, but fared worse on the Ad Fontes dataset, with an accuracy of . This shows that simpler models are not well-suited to the task of ideological estimation, especially when compared to more advanced models.
Zero-shot performance. Figure 1 shows the performance of the three LLMs and baselines across the three datasets. Focusing on the zero-shot setting, we observe that GPT-4o outperforms both Mistral and Llama2, achieving an accuracy of , , and across the YouTube, Ad Fontes, and AllSides datasets. Mistral performs better than Llama2, achieving a accuracy on the YouTube dataset closer to GPT-4o whereas Llama2’s accuracy is comparably low at . On the Ad Fontes dataset, GPT-4o performs better than Mistral at accuracy compared to Mistral’s . Finally, all LLMs perform poorly on the AllSides dataset with GPT-4o performing the best at . Comparing with the baselines, despite training on the full training set unlike LLMs, RoBERTa performed far worse on the Ad Fontes dataset at . In contrast, the MLP performs the worst in the YouTube and Ad Fontes datasets.
Algorithm 1 Improvements. Providing the set of demonstrations selected using the Set-BSR approach to the prompt , Figure 1 also shows how the accuracy of the models improves by increasing the number of demonstrations using Algorithm 1. The trend of GPT-4o outperforming Llama2 and Mistral persists as the number of demonstrations increases. GPT-4o consistently achieves the highest scores across all datasets, suggesting that it is particularly well-suited for ideological estimation tasks. Mistral and GPT-4o showed a steady increase in performance as the number of demonstrations increases, both seeing a and increase in accuracy going from ( to for GPT-4o, and to for Mistral). Llama2 saw increases in accuracy across all datasets from , but grew worse as . We also undertake McNemar’s test for assessing statistical significance of all results in Appendix F.
Including Source and Description increases accuracy but not always. Augmenting the test instance to include source and description increases the performance of the models across all the datasets in the zero-shot setting. Figure 2 shows how model performance is affected by altering to test different combinations of the title, source, and description. We consider the best performing ablation of the baseline MLP and RoBERTa models as described in Appendix B. We observed that the source alone was a strong indicator of ideology in the two news datasets and less so in the YouTube slant dataset. These two datasets comprise news articles from well-known news sources of whom the LLMs are most likely aware of allowing it to predict ideology based on its knowledge of the source’s ideology. The accuracy is poorer in the YouTube slant dataset where the sources can be niche and less popular YouTube channels of whom the LLM is unaware of and performed better when the title and description were provided. For the YouTube dataset, performance for GPT-4o increased by when the source and description was provided along with the title. Llama2 saw an increase of and Mistral saw an increase of when the title and source were provided. When augmenting the title with the source and description, we see significantly higher increases for Ad Fontes, with the performance of GPT-4o increasing by , Llama2 by , and Mistral by . Similar increases are seen for AllSides, where adding just the source improved accuracy. We present and discuss these ablations in more detail in Appendix B, and discuss edge cases where the ideology of the source can be misleading in Appendix E.
6 Conclusion
We explored the efficacy of few-shot learning using LLMs for the task of ideology estimation across diverse textual datasets, including news articles and YouTube videos. Our findings demonstrate that LLMs, particularly GPT-4o, achieve state-of-the-art performance in the few-shot setting, surpassing traditional supervised baselines. By leveraging set-cover based ICL, we addressed challenges of scalability and robustness, showcasing the ability of LLMs to adapt to evolving contexts with minimal supervision.
7 Limitations
Despite the promising results and contributions outlined in this paper, several limitations must be acknowledged to identify areas for future work. The performance of our approach is heavily reliant on the capabilities of pre-trained LLMs such as GPT-4o, Llama2-13B, and Mistral-7B. Variability in model updates or differences in underlying pre-training data could affect reproducibility and consistency across different deployments. In addition, while incorporating metadata, such as source information improved performance, it also introduced the potential for biased predictions. Models may disproportionately rely on known ideological leanings of sources, potentially misclassifying nuanced or neutral content.
Furthermore, our experiments focus on text, and extending these methods to classify ideology in non-textual data such as video or audio remains an open challenge and an important direction for future work. Lastly, we acknowledge that the datasets used for evaluation, while diverse, may not fully capture the breadth of content and ideological nuance present in real-world social media platforms. For instance, the balance of liberal, neutral, and conservative content in the datasets may not reflect natural distributions, and also the majority of content consumed by users is generated by other users themselves, not by news media organizations or YouTube channels, suggesting the need to adapt our approach to user-generated posts and comments.
8 Ethics Statement
Automated ideology estimation carries inherent risks, including misuse for surveillance or censorship. Furthermore, reliance on LLMs can perpetuate biases present in training data, which might influence predictions in unintended ways. While the study evaluates multiple datasets, the generalizability of the approach across all social media platforms—with varying content types and user behaviors–—is yet to be fully established.
Addressing these limitations could involve refining LLM prompting techniques, improving model transparency, exploring alternative datasets, and incorporating mechanisms for bias mitigation and real-time adaptability. Future research should also emphasize ethical frameworks for deploying such systems responsibly, ensuring that advancements in technology contribute positively to societal discourse without exacerbating existing challenges.
Appendix
Appendix A Prompts used for LLMs
Here, we present the prompts that are used for querying the models.
Prompt for querying with just the title.
Classify the following news article titles as ideologically liberal, neutral, or conservative. Titles with no ideological content are classified as neutral. Only respond with the final answer.
Prompt for querying with the title and source.
Classify the following news article titles as ideologically liberal, neutral, or conservative. Titles with no ideological content are classified as neutral. The news source is also specified for additional context. Only respond with the final answer.
Prompt for querying with the title, description, and source.
Classify the following news article titles as ideologically liberal, neutral, or conservative. Titles with no ideological content are classified as neutral. The news description is also specified for additional context. Only respond with the final answer.
Appendix B Ablations on different parameters
In this section, we report on the various ablations of title, source, description, and number of shots for the various baseline and LLM models.
B.1 MLP Baseline
Table 1 shows the accuracy of the MLP model under different title, source, and description ablations. The multi-layer perceptron model achieved good accuracy when the source was provided. In the AllSides news dataset, wherein we observe that the source is highly correlated with article ideology, the model achieves accuracy when just the source is provided. Similarly, MLP achieves accuracy on the YouTube slant dataset when title, source, and article description are provided. Finally, the model achieves a high accuracy of on the Ad Fontes news dataset when just the title and source are provided.
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Desc. | Slant | News | News |
| ✓ | 0.51 | 0.38 | 0.43 | ||
| ✓ | 0.61 | 0.55 | 0.69 | ||
| ✓ | 0.55 | 0.38 | 0.47 | ||
| ✓ | ✓ | 0.63 | 0.62 | 0.57 | |
| ✓ | ✓ | 0.57 | 0.43 | 0.46 | |
| ✓ | ✓ | 0.62 | 0.60 | 0.58 | |
| ✓ | ✓ | ✓ | 0.64 | 0.61 | 0.52 |
B.2 RoBERTa baseline
Table 2 shows the accuracy of the RoBERTa model under different title, source, and description ablations. The RoBERTa model performed well on the YouTube slant dataset achieving an accuracy of when title, source, and description were all provided. The performance was worse on the news datasets, with a maximum of on the Ad Fontes news dataset and on the AllSides news dataset.
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Desc. | Slant | News | News |
| ✓ | 0.57 | 0.35 | 0.52 | ||
| ✓ | 0.45 | 0.48 | 0.48 | ||
| ✓ | 0.62 | 0.35 | 0.50 | ||
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | 0.64 | 0.38 | 0.53 | |
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | ✓ | 0.65 | 0.44 | 0.55 |
B.3 GPT-4o
Table 3 shows the accuracy of the GPT-4o model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best when given a higher number of demonstrations for the YouTube slant and Ad Fontes news datasets. For the YouTube dataset, the model performed best when title, source, and description were all provided whereas for the Ad Fontes news dataset, only title and source were sufficient. The AllSides news dataset showed a different trend in that zero demonstrations gave the best score when only the source was provided. We observed in the test set for this dataset that source and target label were highly correlated (i.e., a video from Fox, a well-known conservative outlet, was almost always labeled conservative) making it simpler for the model to just assign the source ideology as the target ideology. In Appendix E, we investigate this further by considering edge cases in the test set.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.64 | 0.69 | 0.50 | ||
| 0 | ✓ | 0.63 | 0.80 | 0.88 | ||
| 0 | ✓ | 0.56 | 0.64 | 0.44 | ||
| 0 | ✓ | ✓ | 0.64 | 0.78 | 0.74 | |
| 0 | ✓ | ✓ | 0.63 | 0.75 | 0.55 | |
| 0 | ✓ | ✓ | 0.59 | 0.82 | 0.85 | |
| 0 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.72 |
| 4 | ✓ | 0.67 | 0.74 | 0.55 | ||
| 4 | ✓ | ✓ | 0.73 | 0.87 | 0.83 | |
| 4 | ✓ | ✓ | 0.69 | 0.80 | 0.58 | |
| 4 | ✓ | ✓ | ✓ | 0.72 | 0.86 | 0.82 |
| 8 | ✓ | 0.68 | 0.73 | 0.55 | ||
| 8 | ✓ | ✓ | 0.72 | 0.87 | 0.86 | |
| 8 | ✓ | ✓ | 0.70 | 0.81 | 0.59 | |
| 8 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.84 |
| 12 | ✓ | 0.69 | 0.74 | 0.55 | ||
| 12 | ✓ | ✓ | 0.74 | 0.87 | 0.87 | |
| 12 | ✓ | ✓ | 0.71 | 0.82 | 0.59 | |
| 12 | ✓ | ✓ | ✓ | 0.71 | 0.87 | 0.84 |
B.4 Llama2
Table 4 shows the accuracy of the Llama2 model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best at 4 or fewer demonstrations indicating that more demonstrations leads to a drop in performance likely due to a limited context window. For the YouTube and Ad Fontes news datasets, the model performed best when title, source, and only four demonstrations were provided. The AllSides news dataset showed similar trends as GPT-4o in that the source was sufficient without any demonstrations to achieve the best accuracy.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.53 | 0.49 | 0.42 | ||
| 0 | ✓ | 0.61 | 0.52 | 0.82 | ||
| 0 | ✓ | 0.52 | 0.49 | 0.37 | ||
| 0 | ✓ | ✓ | 0.63 | 0.57 | 0.65 | |
| 0 | ✓ | ✓ | 0.58 | 0.57 | 0.46 | |
| 0 | ✓ | ✓ | 0.48 | 0.65 | 0.71 | |
| 0 | ✓ | ✓ | ✓ | 0.58 | 0.69 | 0.59 |
| 4 | ✓ | 0.64 | 0.60 | 0.54 | ||
| 4 | ✓ | ✓ | 0.68 | 0.79 | 0.75 | |
| 4 | ✓ | ✓ | 0.66 | 0.68 | 0.51 | |
| 4 | ✓ | ✓ | ✓ | 0.64 | 0.79 | 0.68 |
| 8 | ✓ | 0.60 | 0.56 | 0.51 | ||
| 8 | ✓ | ✓ | 0.63 | 0.76 | 0.74 | |
| 8 | ✓ | ✓ | 0.63 | 0.67 | 0.52 | |
| 8 | ✓ | ✓ | ✓ | 0.60 | 0.78 | 0.67 |
| 12 | ✓ | 0.60 | 0.60 | 0.50 | ||
| 12 | ✓ | ✓ | 0.62 | 0.75 | 0.72 | |
| 12 | ✓ | ✓ | 0.58 | 0.68 | 0.49 | |
| 12 | ✓ | ✓ | ✓ | 0.56 | 0.78 | 0.67 |
B.5 Mistral
Table 5 shows the accuracy of the Mistral model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best at zero demonstrations indicating that more demonstrations leads to a drop in performance likely due to a limited context window. For the YouTube and Ad Fontes news datasets, the model performed best when title, source, and description were provided with zero demonstrations. The AllSides news dataset again showed similar trends as GPT-4o in that the source was sufficient without any demonstrations to achieve the best accuracy.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.61 | 0.57 | 0.44 | ||
| 0 | ✓ | 0.61 | 0.72 | 0.83 | ||
| 0 | ✓ | 0.52 | 0.59 | 0.34 | ||
| 0 | ✓ | ✓ | 0.67 | 0.72 | 0.71 | |
| 0 | ✓ | ✓ | 0.62 | 0.62 | 0.41 | |
| 0 | ✓ | ✓ | 0.57 | 0.73 | 0.66 | |
| 0 | ✓ | ✓ | ✓ | 0.65 | 0.74 | 0.58 |
| 4 | ✓ | 0.61 | 0.57 | 0.42 | ||
| 4 | ✓ | ✓ | 0.65 | 0.73 | 0.73 | |
| 4 | ✓ | ✓ | 0.62 | 0.61 | 0.44 | |
| 4 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.60 |
| 8 | ✓ | 0.62 | 0.59 | 0.44 | ||
| 8 | ✓ | ✓ | 0.64 | 0.71 | 0.73 | |
| 8 | ✓ | ✓ | 0.63 | 0.62 | 0.48 | |
| 8 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.61 |
| 12 | ✓ | 0.64 | 0.61 | 0.44 | ||
| 12 | ✓ | ✓ | 0.66 | 0.71 | 0.73 | |
| 12 | ✓ | ✓ | 0.61 | 0.63 | 0.48 | |
| 12 | ✓ | ✓ | ✓ | 0.63 | 0.69 | 0.61 |
Appendix C Candidate set size can be smaller
One might argue that the candidate set from which demonstrations are selected is similar to the training set used for training supervised models and, thus, a large amount of annotated data is also required for LLMs. While the size of the candidate set was configured high across all three datasets, we further analyze whether the need for annotated data can be alleviated by considering a smaller candidate set size as well. Specifically, we set the candidate set size to only 500 samples, a fraction of what is generally required to train supervised models. Testing this across all three LLMs and all three datasets in Table 6, we observe that the accuracy is not majorly effected by this change. This is likely because the Set-BSR approach already selects the most optimal demonstrations that maximize coverage.
| LLM | YouTube Slant | Ad Fontes News | AllSides News |
| GPT-4o | 0.66 | 0.74 | 0.52 |
| Llama2 | 0.60 | 0.57 | 0.50 |
| Mistral | 0.63 | 0.59 | 0.42 |
Appendix D Balanced demonstration selection is better than random
We experiment with two demonstration selection techniques namely 1) Random where the demonstrations are randomly selected from the candidate set and 2) Balanced BERTScore where the BERTScore-Recall metric is used to select demonstrations in a label-balanced manner (equal liberal, neutral, and conservative demonstrations). Comparing the performance of Mistral with 8 demonstrations, we see in Table 7 that there is an average increase of in accuracy across the datasets when using Balanced BERTScore selection over random.
| Dataset | Random | Balanced BERTScore |
| YouTube Slant | 0.52 | 0.61 |
| Ad Fontes News | 0.49 | 0.57 |
| AllSides News | 0.39 | 0.42 |
Appendix E Sources can be misleading
Our prior results showed that providing the source information significantly increases the accuracy. However, not all content from a given source may share the ideology of the source itself Haroon et al. (2023), as some liberal articles can come from conservative outlets. To highlight this particular scenario, we identify cases in our test set where the ideology of the article or video is different from the ideology of its source for all three datasets (e.g., a liberal article from Fox News or a conservative article from MSNBC). We rely on AllSides’ media bias chart222https://www.allsides.com/media-bias/media-bias-chart for source ideology and map these sources to their corresponding news articles and YouTube videos. We then compare the accuracy of GPT-4o when 1) only title is provided, 2) only source is provided, and 3) both title and source are provided.
| Liberal Sources | Conservative Sources |
| ABC News | Breitbart |
| Associated Press | CBN |
| CBC News | Daily Caller |
| CBS News | Daily Mail |
| CNN | Daily Wire |
| Daily Beast | Fox Business |
| HuffPost | Fox News |
| Jacobin | New York Post |
| MSNBC | Newsmax |
| Mother Jones | The American Conservative |
| NPR | The American Spectator |
| Slate | The Blaze |
| The Atlantic | The Daily Caller |
| The Guardian | The Epoch Times |
| The Intercept | The Federalist |
| The Nation | The Federalist Society |
| The New York Times | The Post Millennial |
| The New Yorker | The Washington Free Beacon |
| Vox | Washington Examiner |
Table 8 lists the names of the left and right sources identified and mapped to the three datasets. In the YouTube test set, we identified 40 cases where an neutral/conservative article from a liberal source and 21 cases where a liberal/neutral article came from a conservative source. The Ad Fontes news test set had 38 cases of neutral/conservative articles from liberal sources and 11 cases of liberal/neutral articles from conservative sources. Finally, the AllSides news test set had 98 neutral/conservative articles from liberal sources and 0 liberal/neutral articles from conservative sources.
| Non-liberal articles from liberal sources | ||||
| Title | Source | YouTube Slant | Ad Fontes News | AllSides News |
| ✓ | 0.63 | 0.58 | 0.49 | |
| ✓ | 0.47 | 0.38 | 1.00 | |
| ✓ | ✓ | 0.66 | 0.70 | 1.00 |
| Non-conservative articles from conservative sources | ||||
| Title | Source | YouTube Slant | Ad Fontes News | AllSides News |
| ✓ | 0.45 | 0.57 | — | |
| ✓ | 0.00 | 0.00 | — | |
| ✓ | ✓ | 0.00 | 0.52 | — |
Table 9 reports the accuracy of the LLMs on these specific test cases for liberal and conservative sources respectively. We see that the LLMs suffer from incorrect classifications when only the source is considered as the ground truth contradicts the ideology of the source. Here, providing the title increases the accuracy significantly especially in the conservative sources where the accuracy on just the source is for the YouTube and Ad Fontes datasets. In fact, providing the title for the YouTube test set in this case still does not improve the accuracy seeing as the model accuracy is still . This suggests that the LLM overtly relies on the source in this case completely disregarding title wherein accuracy on just the title was close to . Similar trends can be seen for the Ad Fontes news test set where the accuracy of source alone was and title was , and both was . We observe that the AllSides news test set still has better performance when the source is provided at a . On closer inspection, we observe that in the test set sampled, there is a high correlation between the source and article ideology which likely leads to such a high performance when source is provided.
Appendix F Statistical significances
We use the McNemar test to determine the statistical differences between the accuracies of two specific classifiers. Tables 10, 11, and 12 show the significance values for the models trained on the YouTube Slant, Ad Fontes news, and AllSides news datasets respectively.
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 0.13 | 0.06 | 33.06** | 0.05 | 23.58** | 8.77** | 34.18** | 0.21 | 11.05** | 10.75** | 2.56 | 39.32** | 62.01** | 4.50* | 4.21* | 2.34 | 0.39 | 1.46 | 46.89** | 15.42** |
| TS | 0.13 | — | 0.52 | 27.95** | 0.52 | 30.99** | 13.77** | 35.79** | 0.69 | 12.10** | 12.30** | 4.69* | 41.32** | 69.72** | 4.72* | 2.79 | 3.31 | 0.04 | 2.77 | 53.35** | 20.66** | |
| TB | 0.06 | 0.52 | — | 42.29** | 0.00 | 28.96** | 10.27** | 30.30** | 0.06 | 10.25** | 10.16** | 2.13 | 37.80** | 65.91** | 2.96 | 5.33* | 1.70 | 0.88 | 1.07 | 54.30** | 18.06** | |
| TSB | 33.06** | 27.95** | 42.29** | — | 25.95** | 89.19** | 67.10** | 89.28** | 22.45** | 59.74** | 58.72** | 33.34** | 96.75** | 132.08** | 41.44** | 7.87** | 45.75** | 22.90** | 29.36** | 119.71** | 72.90** | |
| S | 0.05 | 0.52 | 0.00 | 25.95** | — | 18.40** | 8.13** | 28.34** | 0.03 | 7.84** | 8.07** | 3.54 | 34.04** | 65.24** | 2.02 | 5.47* | 0.98 | 0.76 | 1.90 | 38.75** | 12.04** | |
| B | 23.58** | 30.99** | 28.96** | 89.19** | 18.40** | — | 4.99* | 3.06 | 13.20** | 0.96 | 0.85 | 5.33* | 5.84* | 20.22** | 5.08* | 42.11** | 11.61** | 29.94** | 6.74** | 7.38** | 0.47 | |
| SB | 8.77** | 13.77** | 10.27** | 67.10** | 8.13** | 4.99* | — | 10.69** | 4.56* | 0.25 | 0.30 | 0.73 | 15.08** | 38.23** | 0.48 | 23.46** | 2.41 | 14.00** | 1.43 | 20.55** | 1.47 | |
| Llama | T | 34.18** | 35.79** | 30.30** | 89.28** | 28.34** | 3.06 | 10.69** | — | 37.09** | 9.03** | 8.62** | 16.12** | 0.35 | 6.28* | 19.19** | 58.08** | 23.16** | 43.91** | 18.64** | 0.02 | 5.61* |
| TS | 0.21 | 0.69 | 0.06 | 22.45** | 0.03 | 13.20** | 4.56* | 37.09** | — | 7.81** | 8.57** | 1.61 | 35.74** | 59.09** | 1.75 | 5.46* | 0.56 | 1.20 | 0.64 | 32.53** | 9.71** | |
| TB | 11.05** | 12.10** | 10.25** | 59.74** | 7.84** | 0.96 | 0.25 | 9.03** | 7.81** | — | 0.00 | 1.69 | 15.84** | 36.84** | 1.71 | 26.06** | 4.14* | 16.75** | 2.70 | 10.61** | 0.16 | |
| TSB | 10.75** | 12.30** | 10.16** | 58.72** | 8.07** | 0.85 | 0.30 | 8.62** | 8.57** | 0.00 | — | 1.92 | 15.09** | 35.27** | 1.83 | 26.38** | 4.38* | 17.32** | 2.98 | 9.67** | 0.11 | |
| S | 2.56 | 4.69* | 2.13 | 33.34** | 3.54 | 5.33* | 0.73 | 16.12** | 1.61 | 1.69 | 1.92 | — | 23.06** | 45.22** | 0.00 | 14.49** | 0.18 | 5.32* | 0.20 | 17.86** | 3.01 | |
| B | 39.32** | 41.32** | 37.80** | 96.75** | 34.04** | 5.84* | 15.08** | 0.35 | 35.74** | 15.84** | 15.09** | 23.06** | — | 5.76* | 21.04** | 69.73** | 31.29** | 53.60** | 25.26** | 0.14 | 10.29** | |
| SB | 62.01** | 69.72** | 65.91** | 132.08** | 65.24** | 20.22** | 38.23** | 6.28* | 59.09** | 36.84** | 35.27** | 45.22** | 5.76* | — | 38.22** | 92.66** | 51.61** | 76.55** | 48.80** | 6.19* | 30.08** | |
| Mistral | T | 4.50* | 4.72* | 2.96 | 41.44** | 2.02 | 5.08* | 0.48 | 19.19** | 1.75 | 1.71 | 1.83 | 0.00 | 21.04** | 38.22** | — | 18.21** | 0.43 | 7.90** | 0.17 | 19.38** | 2.88 |
| TS | 4.21* | 2.79 | 5.33* | 7.87** | 5.47* | 42.11** | 23.46** | 58.08** | 5.46* | 26.06** | 26.38** | 14.49** | 69.73** | 92.66** | 18.21** | — | 13.56** | 2.79 | 11.83** | 69.98** | 32.90** | |
| TB | 2.34 | 3.31 | 1.70 | 45.75** | 0.98 | 11.61** | 2.41 | 23.16** | 0.56 | 4.14* | 4.38* | 0.18 | 31.29** | 51.61** | 0.43 | 13.56** | — | 9.75** | 0.00 | 43.68** | 8.44** | |
| TSB | 0.39 | 0.04 | 0.88 | 22.90** | 0.76 | 29.94** | 14.00** | 43.91** | 1.20 | 16.75** | 17.32** | 5.32* | 53.60** | 76.55** | 7.90** | 2.79 | 9.75** | — | 3.62 | 68.73** | 24.89** | |
| S | 1.46 | 2.77 | 1.07 | 29.36** | 1.90 | 6.74** | 1.43 | 18.64** | 0.64 | 2.70 | 2.98 | 0.20 | 25.26** | 48.80** | 0.17 | 11.83** | 0.00 | 3.62 | — | 20.30** | 4.24* | |
| B | 46.89** | 53.35** | 54.30** | 119.71** | 38.75** | 7.38** | 20.55** | 0.02 | 32.53** | 10.61** | 9.67** | 17.86** | 0.14 | 6.19* | 19.38** | 69.98** | 43.68** | 68.73** | 20.30** | — | 20.33** | |
| SB | 15.42** | 20.66** | 18.06** | 72.90** | 12.04** | 0.47 | 1.47 | 5.61* | 9.71** | 0.16 | 0.11 | 3.01 | 10.29** | 30.08** | 2.88 | 32.90** | 8.44** | 24.89** | 4.24* | 20.33** | — | |
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 34.40** | 20.11** | 449.51** | 341.98** | 188.99** | 377.91** | 96.77** | 32.68** | 38.99** | 224.78** | 66.33** | 47.50** | 172.07** | 51.50** | 2.29 | 18.13** | 288.68** | 232.80** | 135.35** | 287.96** |
| TS | 34.40** | — | 6.67** | 435.86** | 328.99** | 158.38** | 362.25** | 200.31** | 132.25** | 140.91** | 218.99** | 60.72** | 42.21** | 163.93** | 136.48** | 19.91** | 82.16** | 265.40** | 231.29** | 113.12** | 264.55** | |
| TB | 20.11** | 6.67** | — | 440.16** | 344.31** | 173.22** | 374.56** | 163.22** | 81.73** | 103.88** | 233.55** | 68.50** | 50.77** | 176.20** | 104.43** | 3.12 | 62.26** | 283.78** | 240.87** | 128.75** | 283.02** | |
| TSB | 259.15** | 373.77** | 318.02** | — | 24.89** | 175.03** | 15.45** | 68.03** | 125.03** | 124.95** | 107.43** | 249.53** | 286.51** | 146.23** | 130.00** | 286.31** | 175.39** | 83.12** | 75.74** | 204.39** | 74.29** | |
| S | 253.85** | 372.88** | 325.06** | 24.89** | — | 61.13** | 4.88* | 73.52** | 136.62** | 132.00** | 36.33** | 186.29** | 197.09** | 72.16** | 134.29** | 293.99** | 177.19** | 12.26** | 33.39** | 93.04** | 14.84** | |
| B | 199.10** | 272.44** | 239.34** | 175.03** | 61.13** | — | 91.51** | 28.50** | 61.03** | 65.13** | 7.92** | 20.40** | 40.57** | 0.19 | 75.61** | 192.91** | 120.69** | 33.56** | 13.44** | 8.85** | 33.00** | |
| SB | 245.87** | 363.62** | 310.64** | 15.45** | 4.88* | 91.51** | — | 64.08** | 121.96** | 118.26** | 51.81** | 201.48** | 222.84** | 96.82** | 127.00** | 283.93** | 168.95** | 24.57** | 44.52** | 124.66** | 31.57** | |
| Llama | T | 96.77** | 200.31** | 163.22** | 433.77** | 350.42** | 149.02** | 375.46** | — | 21.99** | 21.16** | 236.90** | 72.80** | 54.29** | 180.28** | 18.25** | 138.91** | 40.15** | 276.51** | 263.02** | 114.96** | 266.27** |
| TS | 32.68** | 132.25** | 81.73** | 475.86** | 395.04** | 164.84** | 413.03** | 21.99** | — | 0.00 | 274.00** | 103.51** | 75.98** | 215.70** | 0.00 | 85.23** | 5.80* | 299.13** | 317.34** | 131.29** | 295.40** | |
| TB | 38.99** | 140.91** | 103.88** | 468.49** | 379.28** | 163.60** | 400.96** | 21.16** | 0.00 | — | 271.24** | 99.85** | 76.18** | 220.90** | 0.00 | 81.82** | 8.23** | 299.39** | 307.37** | 133.30** | 297.66** | |
| TSB | 252.24** | 383.38** | 327.86** | 107.43** | 36.33** | 7.92** | 51.81** | 79.93** | 148.70** | 144.11** | — | 78.32** | 130.33** | 12.27** | 136.57** | 313.88** | 184.20** | 7.11** | 2.32 | 28.24** | 5.32* | |
| S | 262.93** | 399.75** | 341.51** | 249.53** | 186.29** | 20.40** | 201.48** | 96.50** | 180.50** | 174.87** | 78.32** | — | 2.49 | 46.08** | 160.35** | 341.84** | 201.12** | 102.45** | 121.28** | 7.80** | 103.54** | |
| B | 260.62** | 393.63** | 345.41** | 286.51** | 197.09** | 40.57** | 222.84** | 107.51** | 190.68** | 185.48** | 130.33** | 2.49 | — | 88.67** | 164.74** | 343.47** | 215.15** | 147.84** | 126.48** | 23.21** | 135.88** | |
| SB | 265.39** | 390.41** | 337.63** | 146.23** | 72.16** | 0.19 | 96.82** | 89.64** | 161.96** | 160.99** | 12.27** | 46.08** | 88.67** | — | 153.62** | 325.90** | 201.51** | 27.24** | 18.31** | 8.15** | 27.36** | |
| Mistral | T | 51.50** | 136.48** | 104.43** | 451.99** | 360.36** | 167.31** | 400.83** | 18.25** | 0.00 | 0.00 | 237.93** | 79.88** | 57.59** | 186.23** | — | 86.05** | 9.88** | 282.52** | 272.88** | 131.00** | 295.80** |
| TS | 2.29 | 19.91** | 3.12 | 465.55** | 366.27** | 173.09** | 401.68** | 138.91** | 85.23** | 81.82** | 249.35** | 84.57** | 60.26** | 194.64** | 86.05** | — | 33.00** | 295.92** | 276.45** | 131.27** | 291.16** | |
| TB | 18.13** | 82.16** | 62.26** | 453.61** | 360.22** | 187.09** | 393.63** | 40.15** | 5.80* | 8.23** | 240.67** | 76.52** | 59.10** | 191.91** | 9.88** | 33.00** | — | 295.82** | 255.13** | 143.73** | 299.24** | |
| TSB | 240.52** | 343.16** | 295.33** | 83.12** | 12.26** | 33.56** | 24.57** | 56.78** | 104.94** | 107.75** | 7.11** | 102.45** | 147.84** | 27.24** | 113.84** | 263.56** | 161.16** | — | 0.71 | 74.25** | 0.02 | |
| S | 270.72** | 411.28** | 350.63** | 75.74** | 33.39** | 13.44** | 44.52** | 97.52** | 180.09** | 169.62** | 2.32 | 121.28** | 126.48** | 18.31** | 163.77** | 345.38** | 205.33** | 0.71 | — | 32.63** | 0.51 | |
| B | 195.26** | 276.06** | 247.61** | 204.39** | 93.04** | 8.85** | 124.66** | 29.48** | 64.62** | 69.14** | 28.24** | 7.80** | 23.21** | 8.15** | 77.65** | 203.68** | 125.25** | 74.25** | 32.63** | — | 77.30** | |
| SB | 219.47** | 317.90** | 275.13** | 74.29** | 14.84** | 33.00** | 31.57** | 47.58** | 94.93** | 95.66** | 5.32* | 103.54** | 135.88** | 27.36** | 104.22** | 243.56** | 146.45** | 0.02 | 0.51 | 77.30** | — | |
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 277.85** | 32.27** | 255.80** | 437.95** | 16.75** | 383.38** | 23.89** | 60.47** | 6.33* | 25.05** | 280.54** | 48.25** | 141.66** | 17.31** | 186.27** | 47.23** | 33.34** | 339.27** | 106.59** | 99.26** |
| TS | 277.85** | — | 195.62** | 3.90* | 139.76** | 316.95** | 94.82** | 258.18** | 26.08** | 212.25** | 81.25** | 28.02** | 301.64** | 3.96* | 272.09** | 4.23* | 357.08** | 139.18** | 62.00** | 448.55** | 43.41** | |
| TB | 32.27** | 195.62** | — | 192.83** | 359.77** | 74.74** | 311.69** | 56.81** | 26.50** | 30.13** | 3.88* | 212.64** | 91.74** | 84.88** | 52.94** | 118.53** | 107.43** | 4.20* | 262.49** | 184.46** | 49.97** | |
| TSB | 255.80** | 3.90* | 192.83** | — | 150.75** | 318.84** | 109.80** | 240.60** | 17.80** | 200.62** | 70.24** | 38.68** | 290.01** | 0.96 | 253.02** | 0.56 | 342.79** | 124.91** | 73.86** | 437.54** | 31.74** | |
| S | 437.95** | 139.76** | 359.77** | 150.75** | — | 518.92** | 27.23** | 483.85** | 187.55** | 434.56** | 284.06** | 20.53** | 540.53** | 150.78** | 481.42** | 149.24** | 547.39** | 324.18** | 57.02** | 656.50** | 263.31** | |
| B | 16.75** | 316.95** | 74.74** | 318.84** | 518.92** | — | 480.17** | 2.13 | 97.45** | 0.63 | 57.42** | 381.03** | 15.78** | 215.47** | 0.13 | 239.73** | 6.63* | 81.30** | 425.04** | 67.27** | 202.21** | |
| SB | 383.38** | 94.82** | 311.69** | 109.80** | 27.23** | 480.17** | — | 441.35** | 150.71** | 393.49** | 243.99** | 7.00** | 508.99** | 119.71** | 437.53** | 109.73** | 505.97** | 279.39** | 21.33** | 619.11** | 229.59** | |
| Llama | T | 23.89** | 258.18** | 56.81** | 240.60** | 483.85** | 2.13 | 441.35** | — | 247.06** | 12.31** | 132.97** | 427.63** | 7.13** | 249.42** | 2.17 | 260.91** | 0.13 | 80.77** | 412.95** | 14.41** | 145.35** |
| TS | 60.47** | 26.08** | 26.50** | 17.80** | 187.55** | 97.45** | 150.71** | 247.06** | — | 164.01** | 24.63** | 106.58** | 224.01** | 12.95** | 162.56** | 17.64** | 151.72** | 16.56** | 119.78** | 203.03** | 0.07 | |
| TB | 6.33* | 212.25** | 30.13** | 200.62** | 434.56** | 0.63 | 393.49** | 12.31** | 164.01** | — | 140.34** | 356.14** | 39.17** | 223.88** | 2.27 | 204.32** | 10.70** | 50.39** | 357.48** | 41.93** | 109.47** | |
| TSB | 25.05** | 81.25** | 3.88* | 70.24** | 284.06** | 57.42** | 243.99** | 132.97** | 24.63** | 140.34** | — | 192.01** | 179.36** | 75.61** | 89.82** | 71.34** | 108.30** | 0.17 | 205.99** | 159.25** | 17.00** | |
| S | 280.54** | 28.02** | 212.64** | 38.68** | 20.53** | 381.03** | 7.00** | 427.63** | 106.58** | 356.14** | 192.01** | — | 510.90** | 72.76** | 365.40** | 51.85** | 441.71** | 207.56** | 1.19 | 541.63** | 131.80** | |
| B | 48.25** | 301.64** | 91.74** | 290.01** | 540.53** | 15.78** | 508.99** | 7.13** | 224.01** | 39.17** | 179.36** | 510.90** | — | 368.67** | 14.31** | 285.29** | 4.71* | 123.77** | 474.60** | 3.24 | 211.43** | |
| SB | 141.66** | 3.96* | 84.88** | 0.96 | 150.78** | 215.47** | 119.71** | 249.42** | 12.95** | 223.88** | 75.61** | 72.76** | 368.67** | — | 214.00** | 0.12 | 281.15** | 69.76** | 93.24** | 372.05** | 15.48** | |
| Mistral | T | 17.31** | 272.09** | 52.94** | 253.02** | 481.42** | 0.13 | 437.53** | 2.17 | 162.56** | 2.27 | 89.82** | 365.40** | 14.31** | 214.00** | — | 292.91** | 4.31* | 87.31** | 396.10** | 30.63** | 140.76** |
| TS | 186.27** | 4.23* | 118.53** | 0.56 | 149.24** | 239.73** | 109.73** | 260.91** | 17.64** | 204.32** | 71.34** | 51.85** | 285.29** | 0.12 | 292.91** | — | 339.67** | 138.56** | 77.25** | 391.42** | 20.98** | |
| TB | 47.23** | 357.08** | 107.43** | 342.79** | 547.39** | 6.63* | 505.97** | 0.13 | 151.72** | 10.70** | 108.30** | 441.71** | 4.71* | 281.15** | 4.31* | 339.67** | — | 175.10** | 462.37** | 26.37** | 227.88** | |
| TSB | 33.34** | 139.18** | 4.20* | 124.91** | 324.18** | 81.30** | 279.39** | 80.77** | 16.56** | 50.39** | 0.17 | 207.56** | 123.77** | 69.76** | 87.31** | 138.56** | 175.10** | — | 232.36** | 226.26** | 35.97** | |
| S | 339.27** | 62.00** | 262.49** | 73.86** | 57.02** | 425.04** | 21.33** | 412.95** | 119.78** | 357.48** | 205.99** | 1.19 | 474.60** | 93.24** | 396.10** | 77.25** | 462.37** | 232.36** | — | 569.29** | 174.80** | |
| B | 106.59** | 448.55** | 184.46** | 437.54** | 656.50** | 67.27** | 619.11** | 14.41** | 203.03** | 41.93** | 159.25** | 541.63** | 3.24 | 372.05** | 30.63** | 391.42** | 26.37** | 226.26** | 569.29** | — | 365.85** | |
| SB | 99.26** | 43.41** | 49.97** | 31.74** | 263.31** | 202.21** | 229.59** | 145.35** | 0.07 | 109.47** | 17.00** | 131.80** | 211.43** | 15.48** | 140.76** | 20.98** | 227.88** | 35.97** | 174.80** | 365.85** | — | |
Appendix G Few-shot results for additional datasets
Figures 3, 4, and 5 show the increases in accuracy across all datasets by increasing the number of shots.
Appendix H Code and Reproducibility
Our code and data used for prompting the models and analyzing the results can be found here: https://anonymous.4open.science/r/llm-ideology-FC41/